OpenStudy (anonymous):

can someone please help me with part B and C?!?!?!?! i will fan and medal! Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (anonymous):

what grade

OpenStudy (anonymous):

alg 1

OpenStudy (anonymous):

What

OpenStudy (anonymous):

its algebra 1

OpenStudy (anonymous):

9th grade

OpenStudy (anonymous):

oh like in high school

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

ok can't help

OpenStudy (anonymous):

why

OpenStudy (anonymous):

I am in 6th

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@YanaSidlinskiy can u help

OpenStudy (yanasidlinskiy):

Do you know what function notation is? Like what is the form of it?

OpenStudy (anonymous):

no thats why i need help

OpenStudy (anonymous):

HELLLOOO

OpenStudy (yanasidlinskiy):

Let (2,225) and (5,480) represent the coordinate points where x= the number of days the boat is rented for; and y= the price of the rental. Given these 2 points, we could write the equation of the straight line. Slope: (480-225)/(5-2)= ....?

OpenStudy (yanasidlinskiy):

Figure that problem out what do you get as your slope?

OpenStudy (anonymous):

I HAVE NO CLUE

OpenStudy (yanasidlinskiy):

Plug that equation into your calculator

OpenStudy (yanasidlinskiy):

Your slope is 85. Now... the equation of the line is: y-225=85(x-2) y-225=85x-170 y=85x+55

OpenStudy (yanasidlinskiy):

So, since we found the slope, we use the same equation using function notation: f(x) = 85x+55

OpenStudy (yanasidlinskiy):

Is it making sense?